car is traveling at a rate of 30 meters per second. What is the car's rate in kilometers per hour? How many kilometers will the car travel in 5 hours? Do not round your answers.
Answer by DrBeeee(684)

This is a tuffy, called unit analysis. You need to convert m/s to km/hr. I'll try to walk you through the steps.
First we have to convert 1 m = (1/1000)km.
Secondly we have 1 h = 3600 s, or 1s = (1/3600)h
Now we take the given rate of 30m/s and multiply by (1/1000) to change m to km, then multiply by 3600 to convert the inverse of s, ie 1/s to 1/h, as follows
(1) 30m/s = 30m/s*(1/1000)km/m*(3600)s/h
Now cancel the units, eg s/s = 1, m/m = 1etc and get
(2) 30m/s = 30*(1/1000)km*(3600)/h or
(3) 30m/s = 30/1000*3600 km/h or
(4) 30m/s = 3*36 km/h or
(5) 30m/s = 108 km/h
The other question of how far the car will travel in 5h is
(6) 5*108km or
(7) 540km.
Answer: 30 k/s is equivalent to 108 km/h. The car will travel 540km in 5 hours.
As I said, it's not easy but very useful for converting units. I hope this helps. Study it a bit, to see what you are doing to convert the units. The hard part is to figure out whether or not you need to muliply or divide by a conversion factor. Do a hundred or so examples and you'll get it. Good luck.